How much water must you add to a 40 L 80% acid solution to reduce it to 20% acid?
Answer by stanbon(75887)

How much water must you add to a 40 L 80% acid solution to reduce it to 20% acid?
----------------
Equation:
water + water = water
0.20*40 + x = 0.80(40+x)
Multiply thru by 100 to get:
------
20*40 + 100x = 80*40 + 80x
20x = 60*40
---
x = 120 L (amt. of water to be added)
=====================================
Cheers,
Stan H.
============